OpenStudy (anonymous):

Please help ill give medals and fan you Solve the compound inequality. 2x – 5 < 3 AND 4 + x ≤ 1

OpenStudy (anonymous):

give me a sec.

OpenStudy (anonymous):

x is less than 4

OpenStudy (anonymous):

x is less than or equal to -3

OpenStudy (anonymous):

@lacrosseplayer Solve the equation. |x + 5| = 2 x = ____ or x = ____

OpenStudy (anonymous):

@lacrosseplayer Which of these equations defines a line that has an y-intercept at (0, 7) and an x-intercept at (-9, 0).

OpenStudy (anonymous):

-3 or -7

OpenStudy (anonymous):

that was for x+5=2

OpenStudy (anonymous):

do you have choices for your second problem? @Lovee_Stasiaa

OpenStudy (anonymous):

7x – 9y = 0 -7x + 9y = 63 7x – 9y = 63 -7x + 9y = 0

OpenStudy (anonymous):

@lacrosseplayer

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

I think its B